OpenStudy (anonymous):

Integrate 7^x * 5^x dx @Calculus1

OpenStudy (ash2326):

7^x*5^x a^x=a^x/ln a 35^x 35^x/ln 35

OpenStudy (anonymous):

I was instructed not to use the property a^x * b^x = (ab)^x.